OpenStudy (anonymous):

I don't understand the first step to eliminating these equations x-y+5z=4 5x+z=0 x+4y+z=-16

OpenStudy (anonymous):

z=-5x x-y-25x=4 x+4y-5x=-16 [-24x-y=4] *4 -4x+4y=-16 ------------- + -100x=0 => x=0 =>z=0 => y=4

OpenStudy (anonymous):

how do you know to do z=-5x

OpenStudy (anonymous):

and how do you solve -5x=z

OpenStudy (anonymous):

5x+z=0 put 5x to the left and you get z=-5x later substitute this value of z in the other equations

OpenStudy (anonymous):

okay and then how does z = 25x

OpenStudy (anonymous):

x-y+5z=4 x-y+5(-5x)=4 x-y-25x=4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so in x-y-25x=4 x+4y+z=-16 what are you solving for